How do I get B in this Equation?
r^2-br+49=0
A=1
B=?
C=49

Answer by mananth(16946)

r^2-br+49=0
When the above equation is a perfect square
(r-7)^2= r^2-2*7*1*r+49
compare the equation r^2-br+49
b=2*7*1
b=14